OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 5x2 + 9y2 = 45.

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

HI again

OpenStudy (misty1212):

first divide by \(45\) to make it in standard form as \[\frac{x^2}{9}+\frac{y^2}{5}=1\]

OpenStudy (misty1212):

is it obvious that the center is \((0,0)\)?

OpenStudy (anonymous):

okay and is the 9=a or 9=a^2?

OpenStudy (anonymous):

and yes, the center is (0,0)

OpenStudy (misty1212):

it is \(a=3\) and since the larger number is under the x, not the y , it looks like the one on the left 

OpenStudy (misty1212):

that means the vertices are 3 units to the LEFT and RIGHT of the center

OpenStudy (misty1212):

again it is important to know what it looks like so the vertices are \((-3,0)\) adn \((3,0)\)

OpenStudy (anonymous):

okay! and then the foci would be (-2,0) and (2,0) right?

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

i.e. yes!

OpenStudy (anonymous):

okay cool thank you! i have two more questions hopefully you can help?

OpenStudy (misty1212):

kk